A bullet is fired vertically into the air from the top of a building and reaches a height of 1600 m above the top of the building. It then falls to the ground below. The total time of flight is 38.0 s.
     (a) Calculate the speed at which the bullet leaves the gun.
     (b) From what height above the ground is the bullet fired?
     (c) Calculate the speed of the bullet when it hits the ground.
 

Grade:12

1 Answers

Himanshu Rathour
5 years ago
Let height of building is h, velocity with which bullet is fired is v
Now
0=u²-2g(1600)
u=177.08m/s
0=v-gt
t=177.08/9.8
≈18s
Total time of flight =38 s
Time of descent =38-18=20s
S=0t+1/2(9.8)20²
S=1960m
Height of building=1960-1600
=360m
Speed with which bullet hits ground is
V=9.8*20=196m/s